OpenStudy (anonymous):

Carl has a piece of toast that has jelly on one side, and he dropped it twice. Both times it landed with the jelly side up. If he drops it two more times, what is the probability that it will have landed jelly side up a total of three times?

OpenStudy (anonymous):

By the way, does anyone viewing have any tips for probability questions because I have another that has to do with a deck of cards. :/

OpenStudy (anonymous):

50% probability for it to have 3 times with the jelly on top and once with the jelly on the bottom?

OpenStudy (anonymous):

a total of 3 times; and it has already landed twice

OpenStudy (anonymous):

its asking what the probability is for it to land a third time

OpenStudy (anonymous):

and only one side of the toast is covered in jelly (1/2)

OpenStudy (anonymous):

ok thanks ^_^

OpenStudy (anonymous):

so if he drops it twice, there's a 1/2 probability for it to land each of time , so 1/2 x 1/2 = 2/4 , get it now?

OpenStudy (anonymous):

Yes I do, thank you

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

He has two more flips, and we want to know the probability that one and only one of those will be jelly side up. Assuming heads/tails 50% probability each flip we have four possible outcomes: U U. U D, D U, D D, and two of the four give us the one Up wanted, so 2/4 = 50% probability.

OpenStudy (anonymous):

@hyntc 1/2 x 1/2 does not equal 2/4, but 1/4 instead

OpenStudy (anonymous):

Yes I agree with douglas